y=sin(2x)/x, x=Pi, dx=.25

find the differentiate


1
Expert's answer
2021-12-16T17:06:19-0500

"\frac{dy}{dx}=\frac{2cos(2x)}{x}-\frac{sin(2x)}{x^2}."

"dy=[\frac{2cos(2x)}{x}-\frac{sin(2x)}{x^2}]dx."

When "x=\pi, \;dx=0.25:"

"dy=[\frac{2cos(2\pi)}{\pi}-\frac{sin(2\pi)}{\pi^2}]*0.25=\frac{0.5}{\pi}\approx 0.1592."